OpenStudy (anonymous):

From a standard deck of 52 cards we randomly select seven without replacement. What is the probability that the 2nd, 3rd and 4th cards are hearts, given that the 1st, 5th, 6th, and 7th cards are hearts?

OpenStudy (anonymous):

Okay, let's start with a simple definition we have: \[ \Pr(A|B) = \frac{\Pr(A\cap B)}{\Pr(B)} \]

OpenStudy (anonymous):

We will say \(A\) is the event that 2,3,4 are hearts. \(B\) is the event that 1,5,6,7 are hearts.

OpenStudy (anonymous):

Can you compute: \(\Pr(A\cap B)\)?

OpenStudy (anonymous):

is it 13/52 *12/51* 11/50* 10/49 *9/48* 8/47* 7/47 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does p(B)= (13/52* 9/48* 8/47* 7/46) + ( 13/52 *12/48 *11/47 *10/46 ) ?

OpenStudy (anonymous):

Do you know how to get \(\Pr(B)\)?

OpenStudy (anonymous):

Getting the probability of \(B\) is a bit tough.

OpenStudy (anonymous):

does my answer right?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

could you help me with that, please?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

First, you have \(13/52\), for the first card.

OpenStudy (anonymous):

For cards 2,3,4 you don't know what will happen, so that much part is tough.

OpenStudy (anonymous):

So you don't know if you'll have 12 cards remaining, or you'll have 9 cards remaining.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

You have 4 cases... Case 1: none of 2,3,4 are hearts Case 2: one of 2,3,4 is heart Case 3: two of 2,3,4 are hearts Case 4: all of 2,3,4 are hearts

OpenStudy (anonymous):

For the first case... There are \( {3 \choose 0} \) ways for it to happen.

OpenStudy (anonymous):

So the probability of \(B\) is given by: \[ \left(\frac{13}{54}\right)\sum_{k=0}^3{3\choose k}\left(\frac{12-k}{48}\right) \left(\frac{11-k}{47}\right) \left(\frac{10-k}{46}\right) \]

OpenStudy (anonymous):

It doesn't look very nice

OpenStudy (anonymous):

13/52 the first card, right

OpenStudy (anonymous):

oh yeah.

OpenStudy (anonymous):

But you understand why it's complicated right?

OpenStudy (anonymous):

yes, i do Thanks I appreciate you help